OpenStudy (nbarrera):

Help+MEDAL! Find a number that is between 3/7 and 5/11

OpenStudy (nbarrera):

\[A. \frac{ 34 }{ 77 } B.\frac{ 7 }{ 14 } C.\frac{ 9 }{ 22 } D. \frac{ 33 }{ 77 }\] These are the answer choices.

OpenStudy (cggurumanjunath):